OpenStudy (anonymous):

Solve the following system of equations. 2x – 3y + z = –1 2x – 2y – 2z = –8 x + 5y – 3z = 2

OpenStudy (anonymous):

ok lets do this

OpenStudy (anonymous):

first of all we want to eliminate a value

OpenStudy (anonymous):

so you choose 2 of the equations and eliminate z

OpenStudy (anonymous):

(2x-3y+z)2 2x-2y-2z

OpenStudy (anonymous):

4x-6y+2z 2x-2y-2z

OpenStudy (anonymous):

Okay, and then what do you do?

OpenStudy (anonymous):

sec i gorgot the consonate

OpenStudy (anonymous):

4x-6y+2x=-2 2x-2y-2z= -8

OpenStudy (anonymous):

6x-8y=-10

OpenStudy (anonymous):

2x – 3y + z = –1 2x – 2y – 2z = –8 x + 5y – 3z = 2

OpenStudy (anonymous):

so now you take the equation we didnt use and eliminate z

OpenStudy (anonymous):

x+5y-3z = 2 (2x-3y+z=-1)3

OpenStudy (anonymous):

x+5y-3z=2 6x-9y+3z=-3

OpenStudy (anonymous):

7x-4y=-3

OpenStudy (anonymous):

7x-4y=-3 6x-8y=-10

OpenStudy (anonymous):

any questions so far?

OpenStudy (anonymous):

those are the 2 values we get after eliminating the z from equations

OpenStudy (anonymous):

so we can solve for x and y now

OpenStudy (anonymous):

(7x-4y=-3)2 6x-8y=-10

OpenStudy (anonymous):

-2 oops

OpenStudy (anonymous):

thank you soo much! i think i understand this a lot better now

OpenStudy (anonymous):

-14x+8y=6 6x-8y=-10

OpenStudy (anonymous):

-8x=-16

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

(7x-4y=-3)6 (6x-8y=-10)-7

OpenStudy (anonymous):

42x-24y=-18 -42x+56y=-70

OpenStudy (anonymous):

opps that +70

OpenStudy (anonymous):

32y=52

OpenStudy (anonymous):

y=52/32

OpenStudy (anonymous):

y=13/8

OpenStudy (anonymous):

so lets check our y value

OpenStudy (anonymous):

6x-8y=-10

OpenStudy (anonymous):

plug in x

OpenStudy (anonymous):

6(2)-8y=-10